# Human-Activity AGV Quality Assessment: A Benchmark Dataset and an Objective Evaluation Metric

### v1

@@ 5.2.1 Quality Scoring Task
Table 3: Ablation study results for different modules across quality metrics. SQA, AQA, and TFE refer to the spatial quality analyzer, action quality analyzer, and text feature extractor, respectively. The results for the three quality scores are reported using SRCC/PLCC, while distortion identification is presented in terms of occurrence and distortion accuracy.
@@
0.633 / 0.638 0.617 / 0.631 0.724 / 0.733 0.575 / 0.512
0.767 / 0.770 0.747 / 0.748 0.728 / 0.740 0.753 / 0.704
0.781 / 0.789 0.754 / 0.765 0.773 / 0.776 0.802 / 0.736
0.812 / 0.821 0.788 / 0.792 0.780 / 0.788 0.815 / 0.752
5.2.2 Body Presence and Distortion Identifying Task
Identifying the presence and distortions of human body parts is a novel task, and we benchmark the performance of video LMM models for this task in Table 2. As for LMMS, GPT-4 series perform best in terms of detecting the presence of body parts, while Qwen2-VL achieve the highest accuracy in identifying whether body parts exhibit semantic distortions. We observe that that identifying distortions in arms is the most challenging, followed by facial distortions while detecting distortions in legs and feet is relatively easier. This aligns with the inherent complexity of structures like faces and arms (particularly hands). Overall, our GHVQ model outperforms in both presence detection and distortion identification tasks, improving accuracy by $1.8\%$ and $28.3\%$, respectively.
5.3 Ablation Study
We conducted an ablation study to validate the effectiveness of the spatial quality analyzer, action quality analyzer, and text feature extractor. The experimental results are shown in Table 3. We observe that the performance across all three quality dimensions and distortion identification degrades the most when the spatial quality analyzer is removed, highlighting the importance of extracting human-focused and AIGC-aware quality features. The action quality analyzer is less crucial than the spatial quality analyzer, possibly because the pre-trained action recognition model is less sensitive to motion in AGVs. The text feature extractor has the least impact compared to the spatial and action quality analyzers, as it only captures video content based on text prompts and does not directly engage with AGVs. In summary, the combination of all three modules achieves the highest performance across all quality dimensions.
6 Conclusion
@@
Supplementary Material
In this supplementary file, we provide more details of Human-AGVQA dataset in Section 7, including text prompts selection, T2V model details, and subjective assessment experiments. We then provide data processing details and further analysis in Section 8, including MOS analysis for T2V models, distortion identification analysis, and the relationship between distortion identification and visual quality. The detailed experiment results are discussed in Section 9, including evaluation criteria in our experiments, compared quality metrics in the proposed benchmark, training details of proposed GHVQ metric, and ablation study of different backbones.
7 Human-AGVQA Dataset
7.1 Text Prompts Selection
Figure 7: The distribution of 44 subcategories in 8 categories of the 400 prompts.
To ensure that the text prompts in our dataset encompass a wide range of real-world human activities, we perform a comprehensive and systematic classification of the words used to create human activity-oriented text prompts. Specifically, we first divide the words into $8$ categories according to their properties: ages, genders, races, occupations, scenes, emotion, appearance, and activities. Each of these categories is then further subdivided into more specific subcategories that are commonly observed in daily human life, as illustrated in Table 4.
Some example text prompts and their corresponding generated videos are illustrated in Figure 9. The proportions of each subcategory within their respective categories are illustrated in Figure 7. It can be observed that the proportions of subcategories in the age, occupation, gender, and ethnicity categories are identical. While for other categories, the subcategories venue and outdoor in scene, sadness in emotion, robust in appearance, and work and socializing in activity have larger proportions than their respective subcategories. In addition, we analyzed the length of text prompts in the Human-AGVQA dataset, and the distribution of word counts are presented in Figure 8. The average text prompt length is $13.85$, with a median of $14$ words.
Figure 8: Histogram and density plot of word count per prompt.
7.2 T2V models
Detailed information about eight T2V methods is presented in Table 5. All selected T2V models have been published within the past two years and represent the most advanced video generation capabilities of T2V models to date. Except for VideoCrafter2, the resolution of videos generated by these models is at least $512$, indicating high clarity. However, the frame rates of StreamingT2V, Latte, and AnimateDiff are below $10$ fps, which may result in low motion consistency in actions.
7.3 Subjective Quality Assessment Experiments
7.3.1 Visual Quality Scoring Criteria
Table 4: Subcategories of ages, genders, races, occupations, scenes, emotion, appearance, and activities.
@@ 7.3.1 Visual Quality Scoring Criteria
Table 5: Video formats generated by the eight T2V models in the Human-AGVQA dataset.
Methods Open Source Year Duration (s) FPS Resolution
Latte [55] 2023 $2.0$ $8$ $512\times 512$
AnimateDiff [22] 2023 $2.0$ $8$ $512\times 512$
Gen-2 [16] 2023 $4.0$ $24$ $1408\times 768$
VideoCrafter2 [9] 2024 $1.6$ $10$ $512\times 320$
MagicTime [98] 2024 $2.1$ $23$ $512\times 512$
StreamingT2V [26] 2024 $5.0$ $8$ $1280\times 720$
Open-sora-plan [40] 2024 $2.7$ $24$ $512\times 512$
StableVideo [1] 2024 $4.0$ $24$ $1024\times 576$
Figure 9: Some example frames of $8$ text categories and $8$ T2V models in proposed Human-AGVQA dataset.
Figure 10: Annotation Interface for Subjective Quality Experiment.
@@
7.3.2 Semantic Distortion Identification Criteria
@@
Figure 11: Annotation Interface for Distortion Identification.
7.3.3 Subjective Quality Experiment Procedure
A total of $80$ subjects participated in the visual quality scoring experiment, with ages ranging from $20$ to $30$ years. The group included $46$ males and $34$ females. Given that the semantic artifact identification task is less complex than the visual quality scoring task, we invited $5$ experts in the field of AIGC quality assessment to perform the semantic artifact identification. Each video was rated by $10$ subjects, and labeled by $5$ experts. The annotation interface for subjective quality experiment and distortion identification are shown in Figure 10 and Figure 11.
All subjects had normal or corrected-to-normal vision. The experiments were conducted in a controlled environment according to the recommendations of ITU-R BT.500-13 [2] to minimize external variables that could influence the judgments of the subjects. The setup included 27-inch calibrated display monitors with 95% DCI-P3 color gamut and a resolution of 4K. The viewing distance was set at $70$ cm. The room lighting was maintained at a consistent level of $300$ lux to ensure uniformity across all viewing sessions.
Before the formal assessments, subjects underwent a training session where they reviewed sample AGVs that were not included in the formal experiment. This session aimed to familiarize them with the evaluation criteria and the rating interface. In the formal experiment, $3,200$ videos were divided into eight groups, each containing $400$ videos that covered all 400 prompts. To avoid visual fatigue, each session lasted no longer than $30$ minutes, ensuring participants could maintain a high level of attention and accuracy in their ratings. In total, there are $96,000$ opinion scores and $160,000$ binary labels in the Human-AGVQA dataset.
8 Data Processing and Analysis
We follow the recommended method in  [2] to process the subjective ratings collected during the experiment. Outlier ratings are detected and removed if they deviate by more than $2\sigma$ (for normal distributions) or $\sqrt{20}\sigma$ (for non-normal distributions) from the mean rating for that condition. Observers contributing more than 5% of outlier ratings are excluded from the analysis. For each test condition, the mean score ($\mu_{i}$) and standard deviation ($\sigma_{i}$) are calculated based on all valid ratings provided by observer $i$:
$\mu_{i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}s_{ij},\quad\sigma_{i}=\sqrt{\frac{1}{N_{i}-1}\sum_{j=1}^{N_{i}}(s_{ij}-\mu_{i})^{2}}$ (9)
@@ 8 Data Processing and Analysis
8.1 Inter-subject Consistency
@@
To evaluate the consistency and reliability of quality scoring, we compute the inter-annotator agreement metric, Krippendorff’s Alpha ($\alpha$), where Krippendorff’s Alpha ($\alpha$) for HA quality, AC quality, OV quality are $0.708$, $0.695$, and $0.628$, respectively, indicating appropriate variations among subjects.
@@
For semantic distortion identification, we calculated the proportions of options receiving 3 votes, 4 votes, and 5 votes, which were $9.7\%$, $41.5\%$, and $46.4\%$, respectively. This demonstrates that the majority of the votes from subjects were consistent.
8.2 MOS Analysis for Each T2V Model
@@
We analyzed the MOS of three quality dimensions across $44$ attributes (mentioned in Table 4) for eight T2V models (mentioned in Table 5), as shown in Figure 12.
Figure 12: The MOS comparison for $8$ text categories of each T2V model.
@@
StableVideo. Overall video (OV) quality emerges as the strongest aspect of StableVideo, consistently achieving high scores on most attributes, including complex scenarios such as workplace and emotion like anger. This highlights the model’s robust capability to generate visually coherent outputs. Human appearance (HA) quality also performs well, though it slightly lags behind overall video quality, indicating the model’s ability to generate clear and recognizable human features but with room for improvement compared to its overall visual performance. In contrast, action continuity (AC) quality represents the model’s primary weakness, consistently scoring the lowest among the three dimensions. This is particularly evident in subcategories requiring complex movements or high dynamism, such as surprise in emotion and sports in activity, revealing the model’s limitations in modeling motion continuity.
AnimateDiff. The AnimateDiff model excels in AC Quality, particularly in emotion and activity, where it consistently outperforms the other quality dimensions. This demonstrates the model’s strength in generating continuous and coherent motion. OV Quality shows stable and relatively high performance across most attributes, reflecting the model’s ability to maintain overall visual consistency. However, HA Quality falls behind, especially in categories such as occupation and activity, where the model struggles with rendering detailed and realistic human appearances. These findings suggest that while AnimateDiff is highly competent in motion continuity, it faces challenges in producing visually intricate details.
Gen-2. The Gen-2 model demonstrates exceptional performance in OV Quality, consistently achieving the highest scores across all subcategories. This underscores the model’s strength in generating visually cohesive outputs. HA Quality ranks second, showing a stable and strong ability to generate clear and recognizable human features, although it slightly trails OV Quality. AC Quality, however, is the weakest of the three dimensions, particularly in subcategories involving complex movements, such as angry in emotion and sports in activity, where it exhibits noticeable limitations.
Latte. Latte shows only average performance in terms of OV quality, but it is still slightly insufficient compared with other models. AC Quality exhibits greater variability, with strong performance in categories like sadness and surprise, but struggles in more complex occupational categories. HA Quality is the weakest dimension, particularly in occupation and activity categories, where the model demonstrates limitations in generating detailed and realistic human appearances. These observations suggest that the Latte model performs moderately well in maintaining overall video quality, its capacity to generate fine-grained human details requires further improvement.
MagicTime. MagicTime performs poorly in AC quality among the three quality scores, and there is a certain gap compared with other models. In addition, HA quality consistently underperforms, especially in categories involving complex emotions, movements, or appearances, highlighting the model’s limitations in rendering detailed human features. The model shows balanced performance in categories like scene and gender but exhibits significant variability in occupation and appearance, indicating its inconsistent generation quality depending on the attribute.
@@
Open-Sora. The performance of OV Quality and AC Quality is not satisfactory, and they still fall short significantly compared to other models. In addition, HA quality consistently lags behind, particularly in the occupation, emotion, and activity categories, where the model struggles to render detailed and realistic human appearances. These results suggest that while Open-Sora excels in visual consistency and motion continuity, it faces challenges in human appearance generation.
StreamingT2V. StreamingT2V shows only average performance in terms of OV quality and HA quality, indicating its limited capability to generate sufficiently detailed human appearances. However, it falls short in AC quality, particularly in dynamic categories such as occupation (Athlete and Cleaner and emotion (surprise and anger), where the scores are significantly lower. These findings indicate that while StreamingT2V is effective in static scenarios, its handling of motion continuity in dynamic contexts requires further improvement.
@@
VideoCrafter2. VideoCrafter2 shows balanced performance in OV quality and AC quality, with stable scores across most attributes. However, HA quality consistently underperforms, particularly in emotion categories such as anger and disgust, and activity categories such as sports and work, where it significantly trails OV and AC quality. These results highlight the model’s limitations in generating detailed and realistic human appearances. Nevertheless, the model demonstrates balanced performance in categories like age, scene, ethnicity, and gender, reflecting its comprehensive generation capability in these areas. However, its inconsistent performance in emotion, particularly in attributes like anger and disgust, underscores the challenges it faces in modeling complex emotional details.
8.3 Distortion Identification Analysis

### v2

@@ 5.2.2. Body Presence and Distortion Identifying Task
Table 3. Ablation study results for different modules across quality metrics. SQA, AQA, and TFE refer to the spatial quality analyzer, action quality analyzer, and text feature extractor, respectively. The results for the three quality scores are reported using SRCC/PLCC, while distortion identification is presented in terms of occurrence and distortion accuracy.
@@
0.625 / 0.631 0.611 / 0.626 0.717 / 0.730 58.36 / 52.16
0.762 / 0.767 0.741 / 0.742 0.721 / 0.734 75.94 / 71.08
0.774 / 0.782 0.752 / 0.761 0.760 / 0.765 81.02 / 73.91
0.805 / 0.809 0.771 / 0.778 0.768 / 0.773 83.02 / 75.94
5.3. Ablation Study
We conducted an ablation study to validate the effectiveness of the spatial quality analyzer, action quality analyzer, and text feature extractor. The experimental results are shown in Table 3.
We observe that the performance across all three quality dimensions and distortion identification degrades the most when the spatial quality analyzer is removed, highlighting the importance of extracting human-focused and AIGC-aware quality features. The action quality analyzer is less crucial than the spatial quality analyzer, possibly because the pre-trained action recognition model is less sensitive to motion in AGVs. The text feature extractor has the least impact compared to the spatial and action quality analyzers, as it only captures video content based on text prompts and does not directly engage with AGVs. In summary, the combination of all three modules achieves the highest performance across all quality dimensions.
6. Conclusion
@@
Appendix A Human-AGVQA Dataset
A.1. Text Prompts Selection
Figure 7. The distribution of 44 subcategories in 8 categories of the 400 prompts.
To ensure that the text prompts in our dataset encompass a wide range of real-world human activities, we perform a comprehensive and systematic classification of the words used to create human activity-oriented text prompts. Specifically, the categories of human activities are derived from the AIGC VQA dataset, Vbench(Huang et al., 2024), FETV(Liu et al., 2023), and real-world video caption datasets, InternVid-10M(Radford et al., 2021a) and Webvid-10M(Bain et al., 2021). We use GPT-4 to identify high-frequency activities within these datasets, and finally we divide the words into $8$ categories according to their properties: ages, genders, races, occupations, scenes, emotion, appearance, and activities. Each of these categories is then further subdivided into more specific subcategories that are commonly observed in daily human life, as illustrated in Table 4. The prompts require at least one person to be present. 60% of the prompts specify a single person, while 40% involve multiple individuals.
The proportions of each subcategory within their respective categories are illustrated in Figure 7. It can be observed that the proportions of subcategories in the age, occupation, gender, and ethnicity categories are identical. While for other categories, the subcategories venue and outdoor in scene, sadness in emotion, robust in appearance, and work and socializing in activity have larger proportions than their respective subcategories. In addition, we analyzed the length of text prompts in the Human-AGVQA dataset, and the distribution of word counts are presented in Figure 8. The average text prompt length is $13.85$, with a median of $14$ words.
Table 4. Subcategories of ages, genders, races, occupations, scenes, emotion, appearance, and activities.
@@ A.1. Text Prompts Selection
Figure 8. Histogram and density plot of word count per prompt.
Table 5. Video formats generated by the $15$ T2V models in the Human-AGVQA dataset.
Methods Duration (s) FPS Resolution Launch Month Utilized Version Version Release Month Open Source
Show-1 (Zhang et al., 2023a) $3.4$ $8$ $1914\times 951$ 2023.10 - 2023.10
VideoCrafter2 (Chen et al., 2024a) $1.6$ $10$ $512\times 320$ 2023.04 - 2024.01
StreamingT2V (Henschel et al., 2024) $5.0$ $8$ $1280\times 720$ 2024.03 - 2024.03
AnimateDiff (Guo et al., 2024) $2.0$ $8$ $512\times 512$ 2024.03 AnimateDiff-Lightning 2024.03
MagicTime (Yuan et al., 2024) $2.1$ $23$ $512\times 512$ 2024.04 - 2024.04
Open-sora-plan (Lab and etc., 2024) $2.7$ $24$ $512\times 512$ 2024.03 v1.1.0 2024.05
Latte (Ma et al., 2024) $2.0$ $8$ $512\times 512$ 2024.01 Latte-1 2024.05
Gen-2 (Esser et al., 2023) $4.0$ $24$ $1408\times 768$ 2023.02 Gen-2 2024.06
StableVideo (AI, 2024a) $4.0$ $24$ $1024\times 576$ 2023.11 - 2024.06
Dreamina (ByteDance, 2024) $5.0$ $24$ $1280\times 951$ 2024.05 S2.0 2024.11
LTX-Video (HaCohen et al., 2024) $4.8$ $25$ $704\times 480$ 2024.11 v0.9.1 2024.12
Ying (AI, 2024b) $5.0$ $30$ $960\times 720$ 2024.07 - 2024.12
Sora (OpenAI, 2024b) $5.0$ $24$ $512\times 512$ 2024.02 480p 2024.12
Kling (Technology, 2024) $4.0$ $24$ $1024\times 576$ 2024.06 kling1.6 2024.12
Wan (Cloud, 2025) $5.3$ $30$ $1280\times 720$ 2023.07 wanx2.1-t2v-turbo 2025.01
A.2. T2V models
Table 5 provides detailed information about 15 T2V models, including 7 commercially available ones. These models are sorted by the release time of the version used. All selected models were released within the past two years, showcasing the state-of-the-art capabilities in video generation for T2V models. The video resolution of most models is at least 512x512, ensuring high clarity. Notably, VideoCrafter2 and LTX-Video generate videos with lower resolutions. In terms of frame rate, some models such as Show-1, StreamingT2V, Latte, and AnimateDiff produce videos with frame rates below 10 fps. This could lead to potential issues with motion consistency in actions, as lower frame rates may cause choppy or less fluid motion. Some example text prompts and their corresponding generated videos are illustrated in Figure 9.
A.3. Subjective Quality Assessment Experiments
A.3.1. Visual Quality Scoring Criteria
@@
Figure 9. Some example frames of $8$ text categories and $8$ T2V models in proposed Human-AGVQA dataset.
@@
A.3.2. Semantic Distortion Identification Criteria
@@
Figure 10. Annotation Interface for Subjective Quality Experiment.
Figure 11. Annotation Interface for Distortion Identification.
A.3.3. Subjective Quality Experiment Procedure
A total of $80$ subjects participated in the visual quality scoring experiment, with ages ranging from $20$ to $30$ years. The group included $46$ males and $34$ females. Given that the semantic artifact identification task is less complex than the visual quality scoring task, we invited $5$ experts in the field of AIGC quality assessment to perform the semantic artifact identification. Each video was rated by $12$ subjects, and labeled by $5$ experts. The annotation interface for subjective quality experiment and distortion identification are shown in Figure 10 and Figure 11.
All subjects had normal or corrected-to-normal vision. The experiments were conducted in a controlled environment according to the recommendations of ITU-R BT.500-13 (Met, 2002) to minimize external variables that could influence the judgments of the subjects. The setup included 27-inch calibrated display monitors with 95% DCI-P3 color gamut and a resolution of 4K. The viewing distance was set at $70$ cm. The room lighting was maintained at a consistent level of $300$ lux to ensure uniformity across all viewing sessions.
Before the formal assessments, subjects underwent a training session where they reviewed sample AGVs that were not included in the formal experiment. This session aimed to familiarize them with the evaluation criteria and the rating interface. In the formal experiment, $6,000$ videos were divided into $15$ groups, each containing $400$ videos that covered all $400$ prompts. To avoid visual fatigue, each session lasted no longer than $30$ minutes, ensuring participants could maintain a high level of attention and accuracy in their ratings. In total, there are $270,000$ opinion scores and $360,000$ binary labels in the Human-AGVQA dataset.
Appendix B Data Processing and Analysis
We follow the recommended method in  (Met, 2002) to process the subjective ratings collected during the experiment. Outlier ratings are detected and removed if they deviate by more than $2\sigma$ (for normal distributions) or $\sqrt{20}\sigma$ (for non-normal distributions) from the mean rating for that condition. Observers contributing more than 5% of outlier ratings are excluded from the analysis. For each test condition, the mean score ($\mu_{i}$) and standard deviation ($\sigma_{i}$) are calculated based on all valid ratings provided by observer $i$:
(9) $\mu_{i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}s_{ij},\quad\sigma_{i}=\sqrt{\frac{1}{N_{i}-1}\sum_{j=1}^{N_{i}}(s_{ij}-\mu_{i})^{2}}$
@@ Appendix B Data Processing and Analysis
B.1. Inter-subject Consistency
@@
To evaluate the consistency and reliability of quality scoring, we compute the inter-annotator agreement metric, Krippendorff’s Alpha ($\alpha$), where Krippendorff’s Alpha ($\alpha$) for HA quality, AC quality, OV quality are $0.643$, $0.658$, and $0.705$, respectively, indicating appropriate variations among subjects.
@@
For semantic distortion identification, we calculated the proportions of options receiving 3 votes, 4 votes, and 5 votes, which were $9.1\%$, $44.6\%$, and $43.8\%$, respectively. This demonstrates that the majority of the votes from subjects were consistent.
B.2. MOS Analysis for Each T2V Model
@@
We analyzed the MOS of three quality dimensions across $44$ attributes (mentioned in Table 4) for $15$ T2V models (mentioned in Table 5), as shown in Figure 12.
Figure 12. The MOS comparison for $8$ text categories of each T2V model.
@@
Sora. Sora excels in all three aspects. Overall video (OV) quality emerges as the strongest aspect of Sora, consistently achieving high scores on most attributes, including complex scenarios such as workplace in scene and emotions like anger. This highlights the robust capacity of the model to generate visually coherent output. Human appearance (HA) quality also performs well. In contrast, action continuity (AC) quality represents the model’s primary weakness in most cases. This is particularly evident in subcategories requiring complex movements or high dynamism, such as anger in emotion and sports in activity, revealing the model’s limitations in modeling motion continuity.
Show1. Show1 demonstrates average performance across the three evaluated dimensions. Notably, the AC Quality is slightly higher than the HA Quality in most cases. Additionally, Show1 maintains a relatively balanced performance across various attribute prompts. This indicates that the model is capable of handling a wide range of prompts with consistent results, without significant bias toward any particular category. However, there are slight performance drops in specific areas, such as the athlete in occupation, disgust in emotion, and sports in activity.
Dreamina. Dreamina performs well across all three evaluation dimensions. Overall, the OV Quality is slightly higher than the other two quality metrics, indicating that the model may face challenges in generating consistent representations of human features and actions. Additionally, Dreamina shows noticeable performance discrepancies across different attribute prompts. For example, there is a clear distinction between the chef and cleaner in occupations, and the model’s representation of fear in emotions and sports in activities is relatively weaker compared to other categories. These variations suggest that it struggles with consistently representing certain categories.
Ying. Ying demonstrates above-average performance across all three evaluation dimensions. In most cases, the AC quality slightly exceeds the other two quality metrics. However, Ying shows noticeable variability in performance across different attribute prompts, particularly in videos involving significant movement. For example, the model’s representation of the athlete in occupation and sports in activity is notably weaker compared to other categories. In contrast, the performance across other attributes shows minimal variation. This suggests that while the model performs well overall, it faces challenges in consistently representing dynamic actions and high-motion scenarios.
LTX LTX shows below-average performance across all three evaluation dimensions. The HA quality is generally lower than the other metrics, suggesting difficulty in accurately depicting human features. Performance varies significantly across attributes, with relatively better results for the middle-aged in age, while other age categories and the painter and student in occupations show poor results. The model also struggles with representing excitement and fear in emotions, as well as sports and study in activities. These discrepancies highlight the model’s challenges in providing consistent and accurate representations across various attributes.
Wan Wan performs exceptionally well across all three evaluation dimensions. In most cases, the OV quality is slightly higher than the other two metrics. Additionally, Wan demonstrates a relatively balanced performance across various attribute prompts, with only slight weaknesses in the athlete in occupation and sports in activity. Performance in other attributes shows no significant differences. This indicates that Wan generally maintains consistent quality, with room for improvement in dynamic or high-motion scenarios.
Kling Kling performs exceptionally well across all three evaluation dimensions. In most cases, similar to Wan, the OV quality is slightly higher than the other two metrics. Kling exhibits a more balanced performance across a wide range of attribute prompts, with only slight weaknesses in the athlete in occupation and sports in activity. Performance in other attributes shows no significant variations, demonstrating strong and consistent quality across all areas. This indicates that Kling delivers reliable results across different prompts.
StableVideo. StableVideo performs above average across all three evaluation dimensions. However, its HA and AC qualities are noticeably lower than OV quality, suggesting challenges in accurately representing human features and maintaining smooth, continuous actions. StableVideo also shows significant variability across different attribute prompts. It performs poorly in the venue in scene, the cleaner in occupation, the disgust in emotion, and sports activity. These discrepancies indicate that the model struggles with generating complex scenes, effectively conveying emotions, and accurately representing dynamic actions.
@@
Open-Sora. The performance of OV Quality and AC Quality of Open-Sora is not satisfactory, and they still fall short significantly compared to other models. In addition, HA quality consistently lags behind, particularly in the occupation, emotion, and activity categories, where the model struggles to render detailed and realistic human appearances. These results suggest that it faces challenges in human appearance generation.
AnimateDiff. AnimateDiff demonstrates moderate performance across all three evaluation dimensions. In most cases, its HA quality is significantly lower than its AC quality. AnimateDiff’s performance in the athlete in occupation and sports in activity is particularly poor, showing clear discrepancies compared to other categories. These findings suggest that while the model performs adequately in certain areas, it struggles with dynamic actions and complex occupations.
@@
VideoCrafter2. VideoCrafter2 shows balanced performance in OV and AC quality, with stable scores across most attributes. However, HA quality consistently underperforms, particularly in emotion such as anger and disgust, and activity such as sports and work, where it significantly trails OV and AC quality. These results highlight the model’s limitations in generating detailed and realistic human appearances.
MagicTime. MagicTime performs poorly in all three qualities, and there is a certain gap compared with other models. In addition, HA quality consistently underperforms, especially in categories involving complex emotions, movements, or appearances, highlighting the model’s limitations in rendering detailed human features. The model exhibits significant variability in occupation and appearance, indicating its inconsistent generation quality depending on the attribute.
Latte. Latte shows below-average performance in all three qualities, but it is still slightly insufficient compared with other models. OV Quality is the weakest dimension, particularly in occupation and activity categories, where the model demonstrates limitations in generating detailed and realistic human appearances. These observations suggest that the Latte model performs moderately well in maintaining overall video quality, and its capacity to generate fine-grained human details requires further improvement.
Gen-2. Gen-2 performs above average across all three evaluation dimensions. However, its AC quality is generally lower than the other metrics, indicating challenges in maintaining smooth actions. The model also shows significant performance variability, particularly in the athlete and cleaner in occupations, as well as the sports in activity, where its performance is notably weaker. This suggests difficulties in accurately representing dynamic actions and certain occupations.
StreamingT2V. StreamingT2V shows below-average performance in all three qualities, indicating its limited capability to generate sufficiently detailed human appearances. It falls short in AC quality, particularly in dynamic categories such as occupation (athlete and cleaner and emotion (surprise and anger), where the scores are significantly lower. These findings indicate that its handling of motion continuity in dynamic contexts requires further improvement.
B.3. Distortion Identification Analysis